Mathematics 328, Spring Semester 2005

Instructor: Andrew Sommese 291 Hurley Phone: 631-6498

email: sommese@nd.edu

Class Website: www.nd.edu/~sommese/math328

Office Hours: Wednesdays 9 to 10:30

Text: Jay L. Devore - “Probability and Statistics for Engineering and the Sciences with Infotrac,” Duxbury Press; 6th edition (January 2004).

Examinations, homework and grades There will be two departmental examinations and one final examination (whose dates and times are listed below). Each departmental exam is a 50 minute exam and will be worth 100 points. The final exam is a two-hour exam and will be worth 150 points. The final exam will cover all the material of the course, with some extra emphasis on the material covered after the second test.

Homework will be assigned regularly, and is an integral part of the course. It will be worth 100 points total. Homework is expected to be turned in on time, and late homework will be penalized (and eventually not accepted if it happens too often).

The total number of possible points for the semester is 450.

Both examinations and homework are conducted under the honor code. While cooperation in doing homework is permitted (and in fact encouraged), copying is not.

A student who misses an examination will receive zero points for that exam unless he or she has written permission from their Dean. Please be aware that travel plans are not considered to be a valid excuse by the Dean.

Exam Dates

Exam 1: Wednesday, February 16, 2005 (in class) Exam 2: Wednesday, March 30, 2005 (in class)

Final:  (time to be announced)  (place to be announced).